# CA-W3D: Leveraging Context-Aware Knowledge for Weakly Supervised Monocular 3D Detection

### v1

@@ Abstract
Weakly supervised monocular 3D detection, while less annotation-intensive, often struggles to capture the global context required for reliable 3D reasoning. Conventional label-efficient methods focus on object-centric features, neglecting contextual semantic relationships that are critical in complex scenes. In this work, we propose a Context-Aware Weak Supervision for Monocular 3D object detection, namely CA-W3D, to address this limitation in a two-stage training paradigm. Specifically, we first introduce a pre-training stage employing Region-wise Object Contrastive Matching (ROCM), which aligns regional object embeddings derived from a trainable monocular 3D encoder and a frozen open-vocabulary 2D visual grounding model. This alignment encourages the monocular encoder to discriminate scene-specific attributes and acquire richer contextual knowledge. In the second stage, we incorporate a pseudo-label training process with a Dual-to-One Distillation (D2OD) mechanism, which effectively transfers contextual priors into the monocular encoder while preserving spatial fidelity and maintaining computational efficiency during inference. Extensive experiments conducted on the public KITTI benchmark demonstrate the effectiveness of our approach, surpassing the SoTA method over all metrics, highlighting the importance of contextual-aware knowledge in weakly-supervised monocular 3D detection.
@@ III-B Dual-to-One Distillation (D2OD)
To ensure effective knowledge distillation, a mean squared error (MSE) loss is employed to minimize the discrepancy between the trainable encoder features and the fused features. During inference, feature concatenation is no longer required, thereby reducing computational overhead. Finally, the trained encoder features are utilized for 3D bounding box regression.
@@ III-C Learning Objectives
$\mathcal{L}_{\text{ray-tracing}}=\begin{cases}\left\|\mathbf{P}-\mathbf{P}_{R}\right\|_{1},&\begin{subarray}{c}\text{if }\text{Ray}_{\mathbf{P}_{\text{cam}}\rightarrow\mathbf{P}}\\ \text{intersects with }b_{3d}\end{subarray},\\[5.0pt] 0,&\text{otherwise.}\end{cases}$ (8)
@@ IV-C Quantatative Analysis
Tab. I compares our work with the benchmark WeakM3D and several fully-supervised methods on KITTI 3D validation set for car category. The results demonstrate that our proposed method consistently outperforms WeaM3D and even some of fully-supervised methods across almost all evaluation metrics and difficulty levels (Easy, Moderate, and Hard). The most significant enhancement is observed for the ap 11 3D metric, where the Easy category exhibits a substantial increase of 3.27. For the ap 11 BEV metric, the proposed approach achieves marginal improvements across all difficulty levels, particularly in the Hard category, highlighting the impact of context-aware weak supervision in monocular 3D detection. The improvements observed under the hard setting can be attributed to the enhanced contextual knowledge model gained through our framework, which equips the model with a stronger capacity to detect co-occurring objects. Consequently, though objects are visually elusive, the model could still identify them.
@@ IV-E Ablation Experiments
Kullback-Leibler Divergence (KL) 42.57 28.00 20.98 35.89 22.01 16.78
@@
Kullback-Leibler Divergence (KL) 16.50 12.59 8.01 10.87 9.09 9.09
@@ V CONCLUSION
In this paper, we introduced CA-W3D, a novel Context-Aware Weak Supervision framework for monocular 3D object detection, designed to address the limitations of conventional weakly supervised methods in capturing global scene context. Our approach employs a two-stage training paradigm: Region-wise Object Contrastive Matching (ROCM), which aligns monocular embeddings with multi-modal knowledge to enhance contextual understanding, and Dual-to-One Distillation (D2OD), which effectively transfers this knowledge to the monocular encoder while preserving spatial fidelity. Extensive evaluations on the KITTI benchmark demonstrate that CA-W3D significantly improves detection performance. Offering a more robust and generalizable solution for real-world autonomous perception.

### v2

@@ Abstract
Weakly supervised monocular 3D detection, while less annotation-intensive, often struggles to capture the global context required for reliable 3D reasoning. Conventional label-efficient methods focus on object-centric features, neglecting contextual semantic relationships that are critical in complex scenes. In this work, we propose a Context-Aware Weak Supervision for Monocular 3D object detection, namely CA-W3D, to address this limitation in a two-stage training paradigm. Specifically, we first introduce a pre-training stage employing Region-wise Object Contrastive Matching (ROCM), which aligns regional object embeddings derived from a trainable monocular 3D encoder and a frozen open-vocabulary 2D visual grounding model. This alignment encourages the monocular encoder to discriminate scene-specific attributes and acquire richer contextual knowledge. In the second stage, we incorporate a pseudo-label training process with a Dual-to-One Distillation (D2OD) mechanism, which effectively transfers contextual priors into the monocular encoder while preserving spatial fidelity and maintaining computational efficiency during inference. Extensive experiments conducted on the public KITTI benchmark demonstrate the effectiveness of our approach, surpassing the SoTA method over all metrics, highlighting the importance of contextual-aware knowledge in weakly-supervised monocular 3D detection. For implementation details: CAW3D
@@ III-B Dual-to-One Distillation (D2OD)
To ensure effective knowledge distillation, a mean squared error (MSE) loss is employed to minimize the discrepancy between the trainable encoder features and the fused features. During inference, the feature concatenation branch is no longer required. Finally, the trained encoder features are utilized for 3D bounding box regression.
@@ III-C Learning Objectives
[…]
@@ IV-C Quantatative Analysis
Tab. I compares our work with the benchmark WeakM3D and several fully-supervised methods on KITTI 3D validation set for car category. The results demonstrate that our proposed method consistently outperforms WeaM3D and even some of fully-supervised methods across almost all evaluation metrics and difficulty levels. The most significant enhancement is observed for the ap 11 3D metric, where the Easy category exhibits a substantial increase of 3.27. For the ap 11 BEV metric, the proposed approach achieves marginal improvements across all difficulty levels, particularly in the Hard category, highlighting the impact of context-aware weak supervision in monocular 3D detection. The improvements observed under the hard setting can be attributed to the enhanced contextual knowledge model gained through our framework, which equips the model with a stronger capacity to detect co-occurring objects. Consequently, though objects are visually elusive, the model could still identify them.
@@ IV-E Ablation Experiments
Kullback-Leibler Divergence (KL)
42.57 28.00 20.98 35.89 22.01 16.78
@@
Kullback-Leibler Divergence (KL)
16.50 12.59 8.01 10.87 9.09 9.09
@@ V CONCLUSION
In this paper, we introduced a novel Context-Aware Weak Supervision framework for monocular 3D object detection to address the limitations of conventional weakly supervised methods in capturing global scene context. We design a two-stage training paradigm: Region-wise Object Contrastive Matching (ROCM), which aligns regional visual and text knowledge to enhance contextual understanding, and Dual-to-One Distillation (D2OD) to effectively transfer this knowledge to the monocular encoder while preserving spatial fidelity. Evaluations on the KITTI benchmark demonstrate a significantly improves detection performance. Offering a more robust and generalizable solution for real-world autonomous perception. However, the current text prompts in Stage 1 (ROCM) are limited to single-word descriptions. For the future work, we will explore more fine-grained and diverse textual expressions to improve regional visual-text alignment.